OpenStudy (anonymous):

Write an equation, in slope intercept form, of the line that passes through the given point and is parallel to (6,1) x-3y=8

OpenStudy (danjs):

x-3y=8 --> y= (1/3)x - (8/3) Slope is 1/3 Line is parallel to this so it has the same slope of 1/3 Point (6,1) slope 1/3 y-1=(1/3)(x-6)

OpenStudy (anonymous):

Thank you